# Learning beyond Teacher: Generalized On-Policy Distillation with Reward Extrapolation

### v1

@@ 2 Related Work
On-Policy Distillation. By sampling trajectories from the student and aligning the student with the teacher’s logit distribution on each token of these student-generated trajectories, on-policy distillation (OPD) (Agarwal et al., 2024; Gu et al., 2024) realizes dense on-policy learning. Empirically, OPD has been shown to achieve faster and more effective distillation than off-policy distillation (Yang et al., 2025a; Lu and Lab, 2025). Recent OPD studies have explored distillation across different model families (Patiño et al., 2025), developed black-box on-policy distillation methods that do not require access to the teacher’s logits (Ye et al., 2025a), and investigated the self-distillation paradigm that leverage the LLM’s in-context capabilities to distill textual context information into its parameters (Yang et al., 2025c; Hübotter et al., 2026; Shenfeld et al., 2026; Zhao et al., 2026).
@@ 3.2 Generalized On-Policy Distillation
First, we re-formulate the OPD objective as
@@ 4.1.1 Experimental Settings
(a) Trends in the average number of tokens and the average accuracy of the on-policy distilled models across different benchmarks under varying reward scaling factors. The teacher for math reasoning tasks is Qwen3-4B-Non-thinking-GRPO-Math, while the teacher for code generation tasks is Qwen3-4B-Non-thinking-GRPO-Code.
@@ 4.1.2 Results of Single-Teacher Distillation
To demonstrate that the improvement of ExOPD over the teacher is not due to less training of the teacher, we compare the evaluation performance of ExOPD and the teacher after an additional 100 steps of RL training. The results in Table 1 show that the teacher with more continued RL training show smaller improvement compared to ExOPD with fewer steps.
@@ 4.1.3 Results of Multi-Teacher Distillation
Table 2: Comparison against off-policy distillation (SFT) and weight extrapolation (ExPO) methods in both single-teacher and multi-teacher settings with same-sized teacher-student pairs. “Teacher” represents the performance of the domain teacher model (Qwen3-4B-Non-thinking-GRPO-Math for math reasoning and Qwen3-4B-Non-thinking-GRPO-Code for code generation), “Student” represents the initial performance of student model Qwen3-4B-Non-Thinking. Each numerical subscript indicates the absolute improvement or degradation compared to the domain teacher model.
@@ Appendix B Detailed Training Settings
Appendix C Prompt Templates

### v2

@@ 2 Related Work
On-Policy Distillation. By sampling trajectories from the student and aligning the student with the teacher’s logit distribution on each token of these student-generated trajectories, on-policy distillation (OPD) (Agarwal et al., 2024; Gu et al., 2024) realizes dense on-policy learning. Empirically, OPD has been shown to achieve faster and more effective distillation than off-policy distillation (Yang et al., 2025a; Lu and Lab, 2025). Recent OPD studies have explored distillation across different model families (Patiño et al., 2025), developed black-box on-policy distillation methods that do not require access to the teacher’s logits (Ye et al., 2025a), and investigated the self-distillation paradigm that leverage the LLM’s in-context capabilities to distill textual context information into its parameters (Yang et al., 2025c; Hübotter et al., 2026; Shenfeld et al., 2026; Zhao et al., 2026; Penaloza et al., 2026).
@@ 3.2 Generalized On-Policy Distillation
First, we re-formulate the OPD objective (Xiao et al., 2025) as
@@ 4.1.1 Experimental Settings
(a) Trends in the average number of tokens and the average accuracy of the on-policy distilled models across different benchmarks under varying reward scaling factors. The teacher for math reasoning tasks is Qwen3-4B-Non-Thinking-RL-Math, while the teacher for code generation tasks is Qwen3-4B-Non-Thinking-RL-Code.
@@ 4.1.2 Results of Single-Teacher Distillation
To demonstrate that the improvement of ExOPD over the teacher is not due to less training of the teacher, we compare the evaluation performance of ExOPD and the teacher after an additional 100 steps of RL training. The results in Table 1 show that the teacher with more continued RL training show smaller improvement compared to ExOPD with fewer steps. We also demonstrate the generalizability and effectiveness of ExOPD when the teacher models are trained with sufficient RL steps (i.e., 1200 RL steps), with the corresponding results provided in Appendix C.
@@ 4.1.3 Results of Multi-Teacher Distillation
Table 2: Comparison against off-policy distillation (SFT) and weight extrapolation (ExPO) methods in both single-teacher and multi-teacher settings with same-sized teacher-student pairs. “Teacher” represents the performance of the domain teacher model (Qwen3-4B-Non-Thinking-RL-Math for math reasoning and Qwen3-4B-Non-Thinking-RL-Code for code generation), “Student” represents the initial performance of student model Qwen3-4B-Non-Thinking. Each numerical subscript indicates the absolute improvement or degradation compared to the domain teacher model.
@@ Appendix B Detailed Training Settings
Appendix C Results of Distillation from Domain Teachers with Sufficient RL Trainings
Here, we show the on-policy distillation results when the domain teachers are trained with sufficient RL steps (i.e., 1200 steps). The experimental settings are the same as that in the Section 4.1. The results in Table 8 demonstrate the generalizability and effectiveness of ExOPD in this case.
Table 8: Distillation results under stronger teachers with sufficient RL trainings. Each numerical subscript indicates the absolute improvement or degradation compared to the domain teacher model.
Method Math Reasoning Code Generation
Teacher 68.2 59.3 37.3 42.9 51.9 88.9 72.5 28.0 63.1
Student 21.5 21.9 10.0 08.0 15.4 74.7 64.7 17.9 52.4
Single-Teacher Distillation
OPD $68.3\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.1}}}$ $58.7\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-0.6}}}$ $\textbf{38.7}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.4}}}$ $41.2\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-1.7}}}$ $51.7\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-0.2}}}$ $89.3\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.4}}}$ $71.3\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-1.2}}}$ $28.0\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.0}}}$ $62.9\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-0.2}}}$
ExOPD $\textbf{68.4}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.2}}}$ $\textbf{59.2}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-0.1}}}$ $38.2\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.9}}}$ $\textbf{42.8}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-0.1}}}$ $\textbf{52.2}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.3}}}$ $\textbf{89.9}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.0}}}$ $\textbf{73.7}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.2}}}$ $\textbf{29.3}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.3}}}$ $\textbf{64.3}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.2}}}$
Multi-Teacher Distillation
OPD $68.2\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.0}}}$ $\textbf{60.2}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.9}}}$ $\textbf{38.5}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.2}}}$ $40.8\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-2.1}}}$ $51.9\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.0}}}$ $86.4\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-2.5}}}$ $72.1\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-0.4}}}$ $27.6\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-0.4}}}$ $62.0\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-1.1}}}$
ExOPD $\textbf{70.1}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.9}}}$ $59.6\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.3}}}$ $37.5\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.2}}}$ $\textbf{42.7}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0.6484375,0.0546875,0.3125}\definecolor[named]{pgfstrokecolor}{rgb}{0.6484375,0.0546875,0.3125}-0.2}}}$ $\textbf{52.5}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.6}}}$ $\textbf{89.5}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+0.6}}}$ $\textbf{73.9}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.4}}}$ $\textbf{29.7}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.7}}}$ $\textbf{64.4}\mathrlap{{}_{\scriptscriptstyle{\color[rgb]{0,0.58984375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.58984375,0}+1.3}}}$
Appendix D Prompt Templates